OpenStudy (anonymous):

Solve: 4x - 3y = -1 3x - 2y = 0

OpenStudy (anonymous):

Just multiply both equations 3*(4x - 3y) = 3*(-1) => 12x -9y = -3 4*(3x - 2y) = 0 => 12x -8y = 0 Now subtract the first from the second: y=3 Then substitute y=3 in one of the previous equations to find x=2

OpenStudy (jiteshmeghwal9):

First try to make one of the variable equal in both of the equation . let's make 'x'. 3(4x-3y)=(-1)3 4(3x-2y)=(0)4 =>12x-9y=-3 12x-8y=0 now, subtract & solve:)

OpenStudy (jiteshmeghwal9):

My method is called elimination method.